assign probability values P(X) to each of the random variable


1
Expert's answer
2021-05-07T11:56:55-0400

 Let X be the outcome of the roll of a die.

Then X is a random variable.

Its possible values are 1, 2, 3, 4, 5, and 6

Each of these possible values has probability

"P(1)=P(2)=P(3)=P(4)=P(5)=P(6)= \dfrac{1}{6}"